# Magnetorheological Characterization of Blood Analogues Seeded with Paramagnetic Particles

### v1

@@ Acknowledgements
S2 Numerical analysis: interparticle distance and chain angle
@@
In Figures S5 and S6 are shown, for $l_{\mathrm{chain}}=4$ and 10 respectively, the instantaneous scaled interparticle distance, $r/d_{\mathrm{p}}$, and the absolute value of the chain angle, $|\phi\rvert$, of the central particle-pair on final simulated cycle.
@@
Figure S7: Oscillatory shear data gathered with the (left) NBa and (right) VBa seeded with 15 wtof MyOne particles under a 50 mT magnetic field. (Top) Pipkin diagrams depicting the Lissajous curves and (bottom) First harmonic loss ($G^{{}^{\prime\prime}}_{1}$, in circular markers) and storage ($G^{{}^{\prime}}_{1}$, in triangular markers) moduli.
@@
Figure S8: Oscillatory shear data gathered with the NBa with 15 wtof MyOne particles under a 50 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). Scaled Chebyshev coefficient spectrum, $e_{\mathrm{n}}/e_{1}$ and $v_{\mathrm{n}}/v_{1}$ (with reconstructed Lissajous curves), variation of the scaled third harmonic coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$) with applied deformation/deformation rate (negative $G_{M}^{\prime}$ values are presented as open markers). Analysed with MITlaos38.
S3.2 5 wtMyOne @ 250 mT (5_MyOne_250)
@@
Figure S9: Oscillatory shear data gathered with the (left) NBa and (right) VBa seeded with 5 wtof MyOne particles under a 250 mT magnetic field. (Top) Pipkin diagrams depicting the Lissajous curves and (bottom) First harmonic loss ($G^{{}^{\prime\prime}}_{1}$, in circular markers) and storage ($G^{{}^{\prime}}_{1}$, in triangular markers) moduli.
@@
Figure S10: Oscillatory shear data gathered with the NBa with 5 wtof MyOne particles under a 250 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). Scaled Chebyshev coefficient spectrum, $e_{\mathrm{n}}/e_{1}$ and $v_{\mathrm{n}}/v_{1}$ (with reconstructed Lissajous curves), variation of the scaled third harmonic coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$) with applied deformation/deformation rate (negative $G_{M}^{\prime}$ values are presented as open markers). Analysed with MITlaos38.
S3.3 15 wtM270 @ 250 mT (15_M270_250)
@@
Figure S11: Oscillatory shear data gathered with the (left) NBa and (right) VBa seeded with 15 wtof M270 particles under a 250 mT magnetic field. (Top) Pipkin diagrams depicting the Lissajous curves and (bottom) First harmonic loss ($G^{{}^{\prime\prime}}_{1}$, in circular markers) and storage ($G^{{}^{\prime}}_{1}$, in triangular markers) moduli.
@@
Figure S12: Oscillatory shear data gathered with the NBa with 15 wtof M270 particles under a 250 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). Scaled Chebyshev coefficient spectrum, $e_{\mathrm{n}}/e_{1}$ and $v_{\mathrm{n}}/v_{1}$ (with reconstructed Lissajous curves), variation of the scaled third harmonic coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$) with applied deformation/deformation rate (negative $G_{M}^{\prime}$ values are presented as open markers). Analysed with MITlaos38.

### v2

@@ Acknowledgements
S2 Bingham model fit to the obtained flow curves
The top row of Figure S5 shows the Bingham model fits to the experimental flow curves. The bottom row displays the dimensionless viscosity curves as a function of the reduced Mason number (Equations 4 and 5 of the main text), along with the Bingham model prediction:
$\tau=\tau_{\gamma}+\eta_{\infty}\dot{\gamma}\Leftrightarrow\frac{\eta}{\eta_{\infty}}=1+\left(\frac{Mn}{Mn^{*}}\right)^{-1}\,$, (S2)
Figure S5: Flow curves with Bingham model fits (top row) and dimensionless viscosity as a function of the reduced Mason number (bottom row), for each tested particle type, concentration and magnetic field density combination, for samples with the NBa (left) and VBa (right) as continuous phase (unseeded data not shown). Secondary flow limit is shown along with the flow curves (top row).
S3 Oscillatory shear measurements with PP50 geometry
[…]
(PP50) Unseeded
Figure S6: Oscillatory shear data gathered with the unseeded (left) NBa and (right) VBa, with a standard, smooth PP50. (Top) Pipkin diagrams depicting the Lissajous curves. The blue and red curves represent, respectively, the elastic (stress vs strain) and viscous (stress vs strain rate) Lissajous curves. The elastic and viscous stresses are also plotted lightly in their respective colours, and the maximum measured stress is presented above each plot. (Bottom) First-harmonic loss ($G^{\prime\prime}_{1}$, in circular markers) and storage ($G^{\prime}_{1}$, in triangular markers) moduli and experimental limits associated with low-torque issues and instrument and sample inertia.
Analysing the Pipkin diagrams, we can now see the expected fluid responses at high deformation ($\gamma_{0}=100$). The NBa returned practically perfectly circular elastic Lissajous curves, and the viscous curve was the expected positive-diagonal. Moreover, the elastic and viscous stress plots also show practically straight lines, indicating the purely viscous linear response we expect from a Newtonian fluid. On the other hand, at this imposed strain, we can see the VBa presenting a slight bending of the viscous Lissajous, an indicator of additional harmonics emerging in the NLVE. Reducing the deformation to $\gamma_{0}=10$, the NBa shows some apparent viscoelastic behaviour at small frequencies, which, looking at the moduli plots, could be due to surface-tension torque. As such, this could also be the cause for the VBa’s viscoelastic response (at $\gamma_{0}=10$ and $\omega<1$ rad/s) despite its moduli being slightly larger. Nevertheless, at this amplitude, increasing the frequency still somewhat returns the NBa’s purely viscous behaviour (despite some oscillations possibly due to inertial issues), while the VBa presents a slight viscoelasticity which agrees with its very subtle elastic nature ($\lambda_{\mathrm{CaBER}}=0.35$ ms), and there is observable bending of the elastic stress which indicates we are still obtaining a non-linear response. Considering lower strains, $\gamma_{0}<10$, because the NBa data is no longer representative of a purely viscous behaviour, we are wary of gathering conclusions of its viscoelastic counterpart. As such, we assume the data for $\gamma_{0}<10$ is no longer representative of either fluid’s rheology.
As discussed, there is evidence of surface tension torque at low frequency up to middling amplitudes (a non-negligible apparent elasticity of the NBa). Looking at the low-torque limit, we can presume, thus, that surface tension has indeed shifted the limit considerably and the data gathered at low frequency and amplitude should be viewed with scepticism. Considering inertial effects, it is challenging to attribute the fault to instrument or sample inertia; however, errors still dominate at high frequencies and small amplitudes. Despite these issues, the increase in diameter, possibly aided by the elimination of the gap-error (the estimated PP50 gap-error is practically insignificantrodrigues2025, $\varepsilon\approx 3$ µm), successfully widened the experimental window, allowing us to visualise the analogues’ response to oscillatory shear. In the current state, this is useful for standard measurements with samples that exhibit subtle rheological properties.
S4 Numerical analysis: spatial discretization
The calculation of the particle dynamics is not directly dependent on the employed mesh, as the only acting forces are inter-particle (dependent on the uniform magnetic field and their own positions) and viscous effects. As such, to effectively model the chain behaviour, the mesh selection criteria should be the quality of the resulting flow field. Our flow, however, is very simple, where a single moving boundary generates a linear velocity profile. Indeed, such a flow field could be effectively modelled by single 1st-order elements along the channel/gap height. However, the lateral open boundaries act as inlets/outlets to large reservoirs where curved streamlines are expected, which ought to require a finer mesh.
We chose to mesh the domain with quadrilateral elements of equal size, which was defined by partitioning the channel/gap height into $n$ segments: $l=h_{\mathrm{c}}/n$, meaning the whole mesh was composed of $6\times n^{2}$ elements. We tested 5 meshes as shown in Figure S7.
Figure S7: Tested meshes, identified by the number of elements along the channel/gap height, with respective number of elements and element size
First, we evaluated the resulting flows in the region of interest, $h_{\mathrm{c}}\leq x\leq h_{\mathrm{c}}$ (dependent on the selected strain amplitude, $\gamma_{0}=1$, which will be discussed further on), by plotting the axial-velocity profiles, $u_{x}$, at two locations: $x=0$ and $x=h_{\mathrm{c}}$ (see Figure 8 and its caption for information on the coordinate system). These profiles, plotted in Figure S8, show that the flow appears to be independent of the selected mesh, with all profiles superimposing in a triangular shape, as expected.
Figure S8: Axial-velocity profiles, $u_{x}$, at $x=0$ and $x=h_{\mathrm{c}}$, for each tested mesh (identified by the number of elements along the channel/gap height, $n$).Data taken at the end of the final oscillatory cycle, $t=3\,(2\pi/\omega)$ ($\gamma_{0}=1$ and $\dot{\gamma}_{0}=0.7$ s-1).
However, as mentioned, differences should appear near the open boundaries where rougher meshes may not be able to accurately represent inlet/outlet flow. Moreover, the velocity profiles along a gap-spanning plane may not be able to convey some subtle differences between the flow fields. To evaluate this, in Figure S9 are shown velocity profiles along the channel at three heights: $y=0.1h_{\mathrm{c}}$, $0.5h_{\mathrm{c}}$ and $0.9h_{\mathrm{c}}$. Indeed, there are discrepancies between the tested meshes, with the roughest mesh, $n=5$, struggling to model the inlet/outlet flow, and increasing $n$ above 20 does not return significant improvements. The entrance effects seem to become negligible at $1.5h_{\mathrm{c}}\gtrsim x\gtrsim 1.5h_{\mathrm{c}}$, meaning the region where the particles will travel should not be affected. Nevertheless, we opted for the $n=20$ mesh, as it seems to yield the best cost/benefit, and it is also similar to other, previously employed configurations rodrigues2025.
Figure S9: Axial-velocity profiles, $u_{x}$, at $y=0.1h_{\mathrm{c}}$, $0.5h_{\mathrm{c}}$ and $0.9h_{\mathrm{c}}$, for each tested mesh (identified by the number of elements along the channel/gap height, $n$). For clarity, only the first half of the channel is shown, $3h_{\mathrm{c}}\leq x\leq 0$ (the flow is symmetric). Data taken at the end of the final oscillatory cycle, $t=3\,(2\pi/\omega)$ ($\gamma_{0}=1$ and $\dot{\gamma}_{0}=0.7$ s-1).
S5 Numerical analysis: interparticle distance and chain angle
@@
In Figures S10 and S11 the instantaneous scaled interparticle distance, $r/d_{\mathrm{p}}$, and the absolute value of the chain angle, $|\phi\rvert$, of the central particle-pair on the final simulated cycle are shown, for $l_{\mathrm{chain}}=4$ and 10, respectively.
@@
Figure S12: Oscillatory shear data gathered with the (left) NBa and (right) VBa seeded with 15 wt% of MyOne particles under a 50 mT magnetic field. (Top) Pipkin diagrams depicting the Lissajous curves and (bottom) First-harmonic loss ($G^{\prime\prime}_{1}$, in circular markers) and storage ($G^{\prime}_{1}$, in triangular markers) moduli.
@@
Figure S13: Oscillatory shear data gathered with the NBa with 15 wt% of MyOne particles under a 50 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). Scaled Chebyshev coefficient spectrum, $e_{\mathrm{n}}/e_{1}$ and $v_{\mathrm{n}}/v_{1}$ (with reconstructed Lissajous curves), variation of the scaled third harmonic coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$) with applied deformation/deformation rate (negative $G_{M}^{\prime}$ values are presented as open markers). Analysed with MITlaosmitlaos.
S6.2 5 wt% MyOne @ 250 mT (5_MyOne_250)
@@
Figure S14: Oscillatory shear data gathered with the (left) NBa and (right) VBa seeded with 5 wt% of MyOne particles under a 250 mT magnetic field. (Top) Pipkin diagrams depicting the Lissajous curves and (bottom) First-harmonic loss ($G^{\prime\prime}_{1}$, in circular markers) and storage ($G^{\prime}_{1}$, in triangular markers) moduli.
@@
Figure S15: Oscillatory shear data gathered with the NBa with 5 wt% of MyOne particles under a 250 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). Scaled Chebyshev coefficient spectrum, $e_{\mathrm{n}}/e_{1}$ and $v_{\mathrm{n}}/v_{1}$ (with reconstructed Lissajous curves), variation of the scaled third harmonic coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$) with applied deformation/deformation rate (negative $G_{M}^{\prime}$ values are presented as open markers). Analysed with MITlaosmitlaos.
S6.3 15 wt% M270 @ 250 mT (15_M270_250)
@@
Figure S16: Oscillatory shear data gathered with the (left) NBa and (right) VBa seeded with 15 wt% of M270 particles under a 250 mT magnetic field. (Top) Pipkin diagrams depicting the Lissajous curves and (bottom) First-harmonic loss ($G^{\prime\prime}_{1}$, in circular markers) and storage ($G^{\prime}_{1}$, in triangular markers) moduli.
@@
Figure S17: Oscillatory shear data gathered with the NBa with 15 wt% of M270 particles under a 250 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). Scaled Chebyshev coefficient spectrum, $e_{\mathrm{n}}/e_{1}$ and $v_{\mathrm{n}}/v_{1}$ (with reconstructed Lissajous curves), variation of the scaled third harmonic coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$) with applied deformation/deformation rate (negative $G_{M}^{\prime}$ values are presented as open markers). Analysed with MITlaosmitlaos.